For f(x)= 1/2x, determine f(1) - f(3)
Answer by algebrahouse.com(1659)

f(x) = (1/2)x

f(1)
= (1/2)(1) {substituted 1 for x}
= 1/2 {multiplied}

f(3)
= (1/2)(3) {substituted 3 for x}
= 3/2 {multiplied}

1/2 - 3/2 {value of f(1) - value of f(3)}
= -2/2 {subtracted numerators over common denominator}
= -1 {reduced}

f(1) - f(3) = -1

For more help from me, visit: www.algebrahouse.com